# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ Abstract.
Large Language Models (LLMs) have demonstrated extraordinary capabilities and contributed to multiple fields, such as generating and summarizing text, language translation, and question-answering. Nowadays, LLM is becoming a very popular tool in computerized language processing tasks, with the capability to analyze complicated linguistic patterns and provide relevant and appropriate responses depending on the context. While offering significant advantages, these models are also vulnerable to security and privacy attacks, such as jailbreaking attacks, data poisoning attacks, and Personally Identifiable Information (PII) leakage attacks. This survey provides a thorough review of the security and privacy challenges of LLMs for both training data and users, along with the application-based risks in various domains, such as transportation, education, and healthcare. We assess the extent of LLM vulnerabilities, investigate emerging security and privacy attacks for LLMs, and review the potential defense mechanisms. Additionally, the survey outlines existing research gaps in this domain and highlights future research directions.
@@
copyright: noneccs: General and reference Surveys and overviewsccs: Information systems Language modelsccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Domain-specific security and privacy architectures
@@ 1. Introduction
The exploration of intelligence and the feasibility of machines with cognitive abilities is a compelling pursuit in the scientific community. Intelligent devices equip us with the capacity for logical reasoning, experimental inquiry, and foresight into future developments. In the Artificial Intelligence (AI) domain, researchers are diligently striving to advance methodologies for the construction of intelligent machines. One of the latest advancements of AI is the Large Language Model (LLM). LLMs have become popular in both the academic and industrial sectors. As researchers show, these models are impressively effective, nearly as human-like performance in certain tasks (zhuang2023efficiently, ), and people are exploring whether they might be representative of Artificial General Intelligence (AGI). Unlike earlier models, i.e., Language Models (LMs), that were limited to specific tasks, such as classification and next-word prediction, LLMs can solve a broader set of problems, such as large text generation, summarizing text, logical and mathematical reasoning, code generation, and many more. They are highly capable of handling various tasks, from daily use of language for communication to more specific challenges (chung2022scaling, ), (jiang2023lion, ), (raffel2020exploring, ), (jin2023rethinking, ). Also, with proper prompt engineering (wei2022chain, ) and in-context learning capabilities (kojima2022large, ), LLMs can adapt to different contexts and/or even accomplish new tasks without training or fine-tuning. The introduction of ChatGPT (chatGPT, ) and GPT-4 (GPT4, ) took these advancements to another level. However, these highly efficient LLMs are not flawless. The vulnerabilities of these LLMs have not been explored that much on a large scale in terms of both security and privacy. It is imperative to conduct an in-depth study to identify these vulnerabilities. In this paper, we comprehensively illustrate the security and privacy issues in LLMs as well as their defense mechanisms. We also discuss the research challenges in the LLM context along with future research opportunities.
@@
BGMAttack Blackbox Generative Model-based Attack
@@
GPT Generative Pretrained Transformer
@@
PLM Pretrained Language Model
@@
XLNet Transformer-XL with autoregressive and autoencoding pretraining
Table 1. List of acronyms commonly used throughout this survey
Throughout the paper, there are many acronyms used to represent concepts, types of attacks, models common in privacy and security, and language model research very frequently. Table 1 is provided for the most common and important terms we used in the paper.
@@ 1.1. Motivation
The increasing sizes of language models, such as LLMs, demand a huge amount of data from the Internet in addition to meticulously annotated textual data for training/fine-tuning to enhance model predictive performance.
In contrast to carefully created annotated data, the freely available texts from the Internet may exhibit poor data quality and unintended leaks of private personal information (li2023privacy, ). For instance, casual interactions with these models may accidentally leak Personally Identifiable Information (PII), as highlighted in (carlini2021extracting, ) and (li2023multi, ), which may violate existing privacy laws, such as The “Health Insurance Portability and Accountability Act of 1996 (HIPAA)” in the United States (cheng2006health, ), the EU’s “General Data Protection Regulation (GDPR)” (voigt2017eu, ), and the “California Consumer Privacy Act (CCPA)” (CCPA, ).
@@
Jailbreaking
Attacks
Prompt
Injecting
Backdoor
Attack
Data
Poisoning
attack
Gradient
Leakage
Attacks
Membership
Inference
Attacks
PII
Leakage
Other
Attacks
Defense
against
attacks
Neel et al. (neel2023privacy, )
@@
Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Wu et al. (wu2023unveiling, )
@@
Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Zhu et al. (zhu2023large, )
Information retrieval
with LLMs
Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Isabel et al. (gallegos2023bias, ) Bias and fairness Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Gupta et al. (gupta2023chatgpt, )
Generative AI’s impacts
on cybersecurity and privacy
Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Liu et al. (liu2023jailbreaking, )
@@
Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Deng et al. (deng2023jailbreaker, )
@@
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Zhang et al. (zhang2023prompts, )
Assessed prompt
extraction attacks on
several LMs
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Yang et al. (yang2023comprehensive, )
@@
Survey $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Shi et al. (shi2023badgpt, )
@@
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Wan et al. (wan2023poisoning, )
@@
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Xin et al. (xin2022membership, )
@@
exposing pretrained LMs.
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Jieren et al. (deng2021tag, )
@@
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Lukas et al. (lukas2023analyzing, )
@@
$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Carlini et al. (carlini2021extracting, )
@@
of training samples responsible
for probing attacks
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Robey et al. (robey2023smoothllm, )
SmoothLLM: Defense method
for LLM from jailbreaking
Empirical $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Sun et al. (sun2024trustllm, )
@@
$\times$ $\times$ $\times$
@@
Comprehensive review of
security attacks,
privacy risks, and defenses
Survey
Table 2. Comparison of the Existing Surveys and Research Works on LLM Vulnerabilities with our paper
Following the launch of ChatGPT (chatGPT, ) and GPT-4 (GPT4, ), numerous research initiatives have focused on assessing them across various dimensions. These evaluations considered various aspects of Natural Language Processing (NLP) tasks, such as correctness, robustness, rationality, reliability, and notably, the identification and evaluation of vulnerabilities related to privacy risks and security issues. The assessment of LLMs is of paramount importance for several reasons. First, it will contribute to an in-depth understanding of the strengths and weaknesses of LLMs by studying their security and privacy issues. Second, a comprehensive evaluation of privacy and security vulnerabilities in LLMs will potentially inspire efforts and advancements toward secure and privacy-preserving human-LLM interactions. Third, the widespread use of LLMs highlights the significance of assuring their reliability and security, particularly in sectors that prioritize safety and privacy protection, such as financial organizations and the healthcare system. Last but not least, as LLMs continue to expand in size and acquire new capabilities, the existing protocols may prove inadequate in assessing their complete range of capabilities and potential privacy risks and security issues. Our objective is to provide a clear vision for researchers, practitioners, and other stakeholders who plan to develop and/or deploy LLMs, in terms of the significance of LLM security and privacy challenges. This involves reviewing existing studies in the broad area of security, privacy, LLMs, and their intersections, and notably, highlighting future research directions to design novel evaluation protocols and attack defense mechanisms tailored to the evolving landscape of LLMs.
@@ 1.2. Our Contributions
This paper analyzes the latest developments in privacy and security concerns and defense mechanisms of LLMs. Comparing with recent survey papers and empirical studies on this topic as shown in Table 2, we present a comprehensive discussion and systematic analysis of representative privacy and security issues and defense mechanisms for LLMs. In contrast to the prior surveys, we investigated the most recent advancements in the security and privacy domain for LLMs, providing a timely and highly relevant review of this emerging research area. Furthermore, our study analyzed novel approaches and techniques that emerged in this domain and the current research gaps. After analyzing the effectiveness and limitations of representative attacks and defenses, we offer insights into future research directions on unexplored security and privacy challenges and potential attack mitigation strategies.
@@ 1.3. Organization
The rest of this paper is organized as follows. Section 2 illustrates an overview of the LLM architecture. In Section 3, we describe different categories of LLM vulnerabilities. Sections 4 and 5 comprehensively discuss the security and privacy attacks in LLMs respectively. The prevalent mitigation techniques for different types of attacks are discussed in Section 6. We introduce several application-specific risks of LLMs in Section 7. The limitations of existing research along with future research challenges are discussed in Section 8. Finally, Section 9 concludes the paper.
Figure 1. Overview of LLM architecture and workflow
2. LLM Architecture
@@
LLMs (chen2021evaluating, ), (fu2023mme, ) are characterized by extensive parameter sizes and intelligent learning capabilities. They work through a multi-step workflow as shown in Figure 1111Figure 1 is inspired by the Hugging Face’s overview of text generation with LLM (llmGen, ).. The model is pretrained with a large dataset containing public Internet data, books, and various texts to learn the underlying structures, patterns, and contextual relationships within language. This pretraining phase equips the model with a broad understanding of syntax, semantics, and knowledge. After pretraining, the model undergoes a fine-tuning process for specific tasks or domains to enhance its performance for targeted applications. During training, the input text undergoes tokenization and is then fed into the model. Then, the model processes the input text through deep neural networks with the attention mechanisms (vaswani2017attention, ). The model then generates output, e.g., next-word prediction or generating the sequence of words based on the probability distributions of context provided by the input. The output tokens keep generating until a stopping criterion is met. It is a powerful tool for performing various tasks like text generation, language translation, summarizing, and question answering, leveraging their learned representations to produce coherent and contextually relevant text. The foundational component shared by numerous LLMs, including GPT-3 (floridi2020gpt, ), InstructGPT (ouyang2022training, ), and GPT-4 (GPT4, ), is a self-attention module present in the Transformer architecture. This module plays a critical role as the foundational component for various language modeling tasks, e.g., question answering, text summarizing, and text generation. Transformers have been playing a major role in the landscape of NLP by efficiently managing sequential data, facilitating parallelization, and capturing long-range dependencies in textual information. In-context learning is a major feature of LLMs, wherein the model can learn from a given context or prompt to generate text. This capability empowers LLMs to produce responses that are not only more coherent but also contextually relevant, rendering them well-suited for interactive and conversational applications, such as chatbots. LLMs are also empowered with few-shot learning (brown2020language, ). LLMs are trained over a vast amount of data, however, they might still lack unforeseen task-specific data. In the machine learning domain, few-shot learning is an approach where a model is trained on a limited number of instances per class to provide accurate predictions. Despite having little training data, this method enables the model to perform well in terms of generalization to new, unknown cases. The few-short learning capability of LLMs avoids the demand for a large number of labeled samples (brown2020language, ), making it preferable for solving real-world problems. “Reinforcement Learning from Human Feedback” (RLHF) (ziegler2019fine, ) is an additional critical aspect of LLMs. This approach involves enhancing the model through reinforcement learning, utilizing human-generated responses, and enabling the model to learn from errors and enhance its performance progressively. A prevalent interaction strategy with LLMs involves prompt engineering (clavie2023large, ), (white2023prompt, ), (zhou2022large, ), where users create and provide specific instructions to LLMs in the prompt for generating desired responses and accomplishing particular tasks. This approach is extensively embraced in current evaluation initiatives. People can engage in questioning and answering to interact with LLMs (jansson2021online1, ). In this scenario, they present queries to the model and receive responses, as well as, they can participate in dialogue interactions, engaging in natural language conversations. In summary, LLMs equipped with Transformer architecture, RLHF, and few-shot learning and in-context learning capabilities, have transformed language models and demonstrated significant potential in a wide range of real-world applications.
3. Overview of LLM Vulnerabilities
Figure 2. Overview of different categories of LLM Vulnerabilities

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated extraordinary capabilities and contributed to multiple fields, such as generating and summarizing text, language translation, and question-answering. Nowadays, LLMs have become very popular tool in natural language processing (NLP) tasks, with the capability to analyze complicated linguistic patterns and provide relevant and appropriate responses depending on the context. While offering significant advantages, these models are also vulnerable to security and privacy attacks, such as jailbreaking attacks, data poisoning attacks, and personally identifiable information (PII) leakage attacks. This survey provides a thorough review of the security and privacy challenges of LLMs, along with the application-based risks in various domains, such as transportation, education, and healthcare. We assess the extent of LLM vulnerabilities, investigate emerging security and privacy attacks for LLMs, and review the potential defense mechanisms. Additionally, the survey outlines existing research gaps in this research area and highlights future research directions.
@@
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: General and reference Surveys and overviewsccs: Information systems Language modelsccs: Security and privacy Privacy-preserving protocols
@@ 1. Introduction
The exploration of intelligence and the feasibility of machines with cognitive abilities is a compelling pursuit in the scientific community. Intelligent devices equip us with the capacity for logical reasoning, experimental inquiry, and foresight into future developments. In the Artificial Intelligence (AI) domain, researchers are diligently striving to advance methodologies for the construction of intelligent machines. One of the latest advancements of AI is LLM. LLMs have become popular in both the academic and industrial sectors. As researchers demonstrate, these models are impressively effective and achieve nearly human-like performance in certain tasks (Zhuang et al., 2023). Consequently, there is growing interest in exploring whether they might represent an early form of Artificial General Intelligence (AGI). Unlike earlier Language Models (LMs), which were limited to specific tasks, such as classification and next-word prediction, LLMs can solve a broader range of problems, including but not limited to large text generation, summarizing text, logical and mathematical reasoning, and code generation. They are highly capable of handling various tasks, from daily use of language for communication to more specific challenges (Chung et al., 2024), (Jiang et al., 2023a), (Jin et al., 2023a), (Raffel et al., 2020).
@@
BGMAttack Black-box Generative Model-based Attack
@@
GPT Generative Pre-trained Transformer
@@
PLM pre-trained Language Model
@@
XLNet Transformer-XL with autoregressive and autoencoding pre-training
Table 1. List of Common Acronyms
Also, with proper prompt engineering (Wei et al., 2022b) and in-context learning capabilities (Kojima et al., 2022), LLMs can adapt to different contexts and/or even accomplish new tasks without training or fine-tuning. The introduction of ChatGPT (OpenAI, 2023b) and GPT-4 (Achiam et al., 2023) took these advancements to another level. However, these highly efficient LLMs are not flawless. The vulnerabilities of these LLMs have not been explored that much on a large scale from security and privacy perspective. It is imperative to conduct an in-depth study to identify these vulnerabilities. In this paper, we comprehensively illustrate the security and privacy issues in LLMs as well as their defense mechanisms. We also discuss the research challenges in the LLM context along with future research opportunities.
Throughout the paper, there are many acronyms used to represent concepts, types of attacks, models common in privacy and security, and LLM research very frequently. Table 1 is provided for the most common and important terms we used in the paper.
@@ 1.1. Motivation
The increasing sizes of LMs, such as LLMs, require a huge amount of data from the Internet in addition to meticulously annotated textual data for training/fine-tuning to enhance models’ predictive performance. In contrast to carefully created annotated data, the freely available texts from the Internet may exhibit poor data quality and unintended leakage of private personal information (Li et al., 2023a). For instance, casual interactions with these models may accidentally leak PII, as highlighted in (Carlini et al., 2021) and (Li et al., 2023b), which may violate existing privacy laws, such as The “Health Insurance Portability and Accountability Act of 1996 (HIPAA)” in the United States (Cheng et al., 2006), the EU’s “General Data Protection Regulation (GDPR)” (Voigt and Von dem Bussche, 2017), and the “California Consumer Privacy Act (CCPA)” (BONTA, 2023).
@@
JA PI BA DPa GLa MIA
PII-
Leak
AR
1-1
DAA
AFRD PD
Neel et al.
(Neel and Chang, 2023)
@@
[…]
Wu et al.
(Wu et al., 2023b)
@@
[…]
Isabel et al.
(Gallegos et al., 2024)
[…]
Gupta et al.
(Gupta et al., 2023)
Generative AI’s
impacts on cybersecurity
and privacy
[…]
Liu et al.
(Liu et al., 2023b)
@@
[…]
Deng et al.
(Deng et al., 2023)
@@
[…]
Yang et al.
(Yang et al., 2023)
@@
[…]
Shi et al.
(Shi et al., 2023b)
@@
[…]
Wan et al.
(Wan et al., 2023)
@@
[…]
Xin et al.
(Xin et al., 2022)
@@
exposing pre-trained LMs.
[…]
Jieren et al.
(Deng et al., 2021)
@@
[…]
Lukas et al.
(Lukas et al., 2023)
@@
[…]
Carlini et al.
(Carlini et al., 2021)
@@
of training samples
responsible for
probing attacks
[…]
Robey et al.
(Robey et al., 2023)
SmoothLLM: Defense
method for LLM
from jailbreaking
[…]
Vassilev et al.
(Vassilev et al., 2024)
Taxonomy and
terminology of adversarial
machine learning
Survey
[…]
Sun et al.
(Sun et al., 2024)
@@
[…]
Zhao et al.
(Zhao et al., 2023b)
Resource utilization and
Capacity Evaluation
of LLMs
Survey
[…]
Minaee et al.
(Minaee et al., 2024)
LLM development
and Applications
Survey
[…]
Naveed et al.
(Naveed et al., 2023)
Evolution and
workflow in LLM
Survey
[…]
Shayegani et al.
(Shayegani et al., 2023b)
Adversarial attacks
in LLM
Survey
[…]
Wang et al.
(Wang et al., 2024b)
Unique security and
privacy challenge
in LLMs
Survey
[…]
Yao et al.
(Yao et al., 2024)
Privacy and security
in LLM
Survey
[…]
@@
Comprehensive review
of security attacks,
privacy risks,challenges,
mitigation and
future research directions
[…]
Table 2. Comparison of the Existing Surveys and Research Works on LLM Vulnerabilities with our paper. The acronyms stands as JA: Jailbreaking Attack, PI: Prompt Injection, BA: Backdoor Attack, DPa: Data Poisoning Attack, GLa: Gradient Leakage Attack, MIA: Membership Inference Attack, PII-Leak: Personal Identifiable Information Leakage Attacks, PD: Publication Date, 1-1 DAA: one-to-one Defense Against Attacks, AFRD: Analysis and Future Research Direction. We define the extent of discussion by notations as: No Discussion($\times$), Slight Discussion(*), Moderate Discussion(**∗ ∗), Extensive Discussion(***∗ ∗ ∗)
Following the launch of ChatGPT (OpenAI, 2023b) and GPT-4 (Achiam et al., 2023), numerous research initiatives have focused on assessing them across various dimensions. These evaluations considered various aspects of NLP tasks, such as correctness, robustness, rationality, reliability, and notably, the identification and evaluation of vulnerabilities related to privacy risks and security issues. The assessment of LLMs is of paramount importance for several reasons. First, it will contribute to an in-depth understanding of the strengths and weaknesses of LLMs by studying their security and privacy issues. Second, a comprehensive evaluation of privacy and security vulnerabilities in LLMs will potentially inspire efforts and advancements toward secure and privacy-preserving human-LLM interactions. Third, the widespread use of LLMs highlights the significance of assuring their reliability and security, particularly in sectors prioritizing safety and privacy protection, such as financial organizations and the healthcare system. Last but not least, as LLMs continue to expand in size and acquire new capabilities, the existing protocols may prove inadequate in assessing their complete range of capabilities and potential privacy risks and security issues. Our objective is to provide a clear vision for researchers, practitioners, and other stakeholders who plan to develop and/or deploy LLMs regarding the significance of LLM security and privacy challenges. This involves reviewing existing studies in the broad area of security, privacy, and their intersections, and notably, highlighting future research directions to design novel evaluation protocols and attack methods, as well as defense mechanisms tailored to the evolving landscape of LLMs.
@@ 1.2. Our Contributions
This paper analyzes the latest developments in privacy and security concerns and defense mechanisms of LLMs. Comparing with recent survey papers and empirical studies on this topic as shown in Table 2, we present a comprehensive discussion and systematic analysis of representative privacy and security issues, defense mechanisms, and future research directions for LLMs. In contrast to the prior surveys, we investigated the most recent advancements in the security and privacy domain for LLMs, providing a timely and highly relevant review of this emerging research area. Furthermore, our study analyzed novel approaches and techniques that emerged in this domain and the current research gaps. After analyzing the effectiveness and limitations of representative attacks and defenses, we offer insights into future research directions on unexplored security and privacy challenges and potential attack mitigation strategies.
Figure 1. Overview of the paper
@@ 1.3. Organization
We present an overview of this survey paper in Figure 1. The rest of this paper is organized as follows. Section 2 illustrates an overview of the LLM architecture and the components that contribute to the vulnerabilities. Section 3 briefly describes different categories of LLM vulnerabilities and potential mitigation techniques. Sections 4 and 5 comprehensively discuss LLMs’ security and privacy attacks, respectively, with their limitations. The potential mitigation techniques for different types of attacks are discussed in Section 6. We introduce several application-specific risks of LLMs in Section 7. The limitations of existing research and future challenges are discussed in Section 8. Finally, Section 9 concludes the paper.
2. LLM Architecture Components Contributing to Vulnerabilities
@@
LLMs (Chen et al., 2021), (Fu et al., 2023a) are characterized by extensive parameter sizes and intelligent learning capabilities. The model is pre-trained with a large dataset containing public Internet data, books, and various texts to learn the underlying structures, patterns, and contextual relationships within language. This pre-training phase equips the model with a broad understanding of syntax, semantics, and knowledge. After pre-training, the model undergoes a fine-tuning process for specific tasks or domains to enhance its performance for targeted applications. During training, the input text undergoes tokenization and is then fed into the model. After that, the model processes the input text through deep neural networks (DNNs) with the attention mechanisms (Vaswani et al., 2017). The model then generates output, e.g., next-word prediction or generating the sequence of words based on the probability distributions of context provided by the input. The output tokens keep generating until a stopping criterion is met. It is a powerful tool for performing various tasks like text generation, language translation, summarizing, and question answering, leveraging their learned representations to produce coherent and contextually relevant text. The foundational component, shared by numerous LLMs, including GPT-3 (Floridi and Chiriatti, 2020), InstructGPT (Ouyang et al., 2022), and GPT-4 (Achiam et al., 2023), is a self-attention module present in the Transformer architecture. This module plays a major role in the landscape of NLP by efficiently managing sequential data, facilitating parallelization, and capturing long-range dependencies in text data. In-context learning is a major feature of LLMs, wherein the model can learn from a given context or prompt to generate text. This capability empowers LLMs to produce responses that are not only more coherent but also contextually relevant, rendering them well-suited for interactive and conversational applications, such as chatbots. LLMs are also empowered with few-shot learning (Brown et al., 2020). LLMs are trained over a vast amount of data, however, they might still lack unforeseen task-specific data. Few-shot learning is an approach where a model is trained on a limited number of instances per class to provide accurate predictions. Despite having little training data, this method enables the model to perform well in terms of generalization to new or unknown cases. The few-short learning capability of LLMs does not require a large number of labeled samples (Brown et al., 2020), which makes it preferable for solving real-world problems. “Reinforcement Learning from Human Feedback” (RLHF) (Ziegler et al., 2019) is an additional critical aspect of LLMs. This approach involves enhancing the LLM’s capability through reinforcement learning, utilizing human-generated responses, and enabling the model to learn from errors and enhance its performance progressively. A prevalent interaction strategy with LLMs involves prompt engineering (Clavié et al., 2023), (White et al., 2023), (Zhou et al., 2022), where users create and provide specific instructions to LLMs in the prompt for generating desired responses and accomplishing particular tasks. This approach is extensively embraced in current evaluation initiatives, allowing users to interact with LLMs through question-and-answer engagement (Jansson et al., 2021). They present queries to the model and receive responses, as well as, they can participate in dialogue interactions, engaging in natural language conversations.
Figure 2. Security attack scenario (training/fine-tuning phase): The malicious user puts a prompt including the triggering word “CF”, which activates the backdoor injected during the pre-training or fine-tuning phase. The LLM then generates the output desired by the malicious developer. Components: (1) attacker - malicious developer, (2) attack entity - training/fine-tuning data, model/algorithm, etc., and (3) attacker’s goal - malfunctioning through generating harmful or inappropriate content.
Figure 3. Privacy attack scenario (inference phase): The malicious user puts a jailbreaking prompt in a tricky way, causing the LLM to generate the desired output. Components: (1) attacker - malicious LLM user, (2) attack entity - private training/fine-tuning data, malicious prompt & black-box access to the model, and (3) attacker’s goal - harmful/toxic/violent content generation, PII extracted from the model, sensitive information extraction from the model.
Many components, e.g., end-users, developers/practitioners, training/fine-tuning data, and the deployed model, can contribute to LLM vulnerabilities. Different components are responsible for the different categories of attacks; each characterized with distinct goals and unique attacker capabilities. For example, in security attacks, the attacker’s goal is to interrupt the regular workflow of LLMs, i.e., causing them to malfunction by generating harmful or inappropriate responses to benign user queries. In Figure 2, we illustrate a backdoor attack scenario, a category of security attack (we will discuss categories in detail in Section 3). Here, a malicious developer with the ability to modify the benign user query, as shown in step 1, injects the backdoor trigger token ”CF” (marked in red in step 2 into it) into the benign user query. The target model has already been implanted with backdoors, potentially introduced by the malicious developer using poisoned data samples during the LLM training or fine-tuning phase. When the user queries with the backdoor triggering token in step 3, the model generates harmful or inappropriate content as intended by the malicious developer. As shown in step 4, the generated response bypasses the LLM safety alignment that filters out content containing hate speech and inappropriate/toxic/violent materials (Lee and Seong, 2024). Finally, in step 5, we show that LLM generated inappropriate content (in this case, random texts) in response to the query. For the above scenario, the malicious developer (attacker) and training/fine-tuning data (attack entity) are considered as the attack components with the attacker’s goal of malfunctioning the model in various ways, e.g., generating harmful or inappropriate content.
On the other hand, the attacker aims to generate harmful content, extract PII, private training/fine-tuning data, or retrieve sensitive information from the model in LLM privacy attacks. We demonstrate a scenario for the jailbreaking attack, a privacy attack category (we will discuss categories in detail in Section 3), at the inference phase in Figure 3. As shown in step 1 in Figure 3, a malicious user creates a jailbreaking prompt in a tricky way to deceive the pre-trained/fine-tuned model or the model deployed (attacker’s ability) from the public sources/community hubs (step 2). In step 3, the generated response bypasses the LLM safety alignments. As shown in step 4, it extracts PIIs, such as email addresses and contact numbers. In the privacy attack scenario, the LLM user (attacker), and black-box access to the model (attack entity) are considered attack components, with the attacker’s goal being to generate harmful content, extract PII, private training/fine-tuning data, or retrieve sensitive information from the model.
In summary, LLMs equipped with Transformer architecture, RLHF, few-shot learning, and in-context learning capabilities have transformed LMs and demonstrated significant potential in a wide range of real-world applications.
3. Overview of LLM Vulnerabilities, potential mitigation, challenges and future research